# Estimating non-linear functionals of trawl processes

### v2

@@ 4.2 Proof of Theorems 1 - 4
Without loss of generality we may and do assume that $g\geq 0$ (otherwise analyze $g^{+}$ and $g^{-}$ separately). Thus, we only need to check that for every $0\leq t\leq T$, $\Psi_{t}^{n}(g)\overset{\mathbb{P}}{\rightarrow}\Psi_{t}(g)$. Suppose first that $g$ is bounded. Then
@@
Theorem 1 in [17] along with the boundedness of $g$ allow us to apply the Dominated Convergence twice to conclude that the right-hand side of the previous inequality goes to 0 as $n\rightarrow\infty$. Now suppose that $g$ is of polynomial growth of order $q>0$ (when $q=0$ $g$ is bounded). Let $\psi:\mathbb{R}^{+}\rightarrow[0,1]$ be a function of class $\mathcal{C}^{\infty}$ such that $\mathbf{1}_{[1,\infty)}\leq\psi\leq\mathbf{1}_{[1/2,\infty)}$, and for every $m\in\mathbb{N}$, set $\psi_{m}(x):=1-\psi(\rvert x\rvert/m)$ as well as $\tilde{\psi}_{m}(x)=\psi(\rvert x\rvert/m).$ With this notation, we decompose
@@
where $g_{m}=g\psi_{m}$ and $\tilde{g}_{m}=g\tilde{\psi}_{m}$. Since for each $m$, $g_{m}$ is continuous and bounded, the first part of the proof gives that $\Psi_{t}^{n}(g_{m})\overset{\mathbb{P}}{\rightarrow}\int_{0}^{t}g_{m}(a(s))\mathrm{d}s$. In view that $\int_{0}^{t}g_{m}(a(s))\mathrm{d}s\rightarrow\Psi_{t}(g),$ as $m\rightarrow\infty$, we are left to show that for every $\eta>0$
@@
Since $g$ is of polynomial growth of order $q>0$, we have that for every $m\geq 1$, $\rvert\tilde{g}_{m}(x)\rvert\leq C_{q}\rvert x\rvert^{q}\mathbf{1}_{\rvert x\rvert>m}$, by the definition of $\psi$. Therefore, (43) will be achieved if
@@
Relation (25) implies that
$Y_{n}^{m}(t)$ $\lesssim_{q}\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert\tilde{a}(l\Delta_{n})\rvert^{q}\mathbf{1}_{\{\rvert\hat{a}(l\Delta_{n})\rvert>m\}}$
$+\rvert\frac{1}{n}S_{n}\rvert^{q}\frac{1}{(n\Delta_{n})^{p}}\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{q}$
$=:C_{q}(Y_{n}^{m,1}+Y_{n}^{2}).$
As in the proof of Lemma 4, we see that $Y_{n}^{2}\overset{\mathbb{P}}{\rightarrow}0$. Hence,
@@
$\lesssim$ $\frac{t}{(n\Delta_{n})^{q/2}}+\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}a(s)^{q}\mathrm{d}s\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m).$
By Markov’s and Hölder’s inequalities, we further have that
$\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m)\leq$ $\mathbb{P}(\rvert\tilde{a}(l\Delta_{n})\rvert>m/2)+\mathbb{P}(\rvert\frac{1}{n\Delta_{n}}(X_{n\Delta_{n}}-X_{l\Delta_{n}})\frac{1}{n}S_{n}\rvert>m/2)$
$\lesssim$ $\frac{1}{m}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})\rvert)+\frac{1}{m}\frac{1}{n\Delta_{n}}\mathbb{E}(\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{2})^{1/2}\mathbb{E}(\rvert\frac{1}{n}S_{n}\rvert^{2})^{1/2}.$
Therefore,
$\mathbb{E}(\rvert Y_{n}^{m,1}\rvert)\lesssim\frac{1}{(n\Delta_{n})^{q/2}}+\frac{1}{m}\left(1+\frac{1}{n\Delta_{n}}\mathbb{E}(\rvert\frac{1}{n}S_{n}\rvert^{2})^{1/2}\right).$
@@
Our next goal is to unify the proof of Theorems 2 and 4. To do so, we use $\mathcal{S}^{n}(g)$ ($\mathcal{S}(g)$) to denote either $\Psi^{n}(g)$, $\Lambda^{n}(g)$, or $\bar{\Lambda}^{n}(g)$ ($\Psi(g)$, $\Lambda(g)$, or $\bar{\Lambda}(g)$) and note that
@@
Then, if $g$ is of class $\mathcal{C}^{1}$, we deduce from the Mean-Value Theorem that there is $\hat{a}(l\Delta_{n})\land a(l\Delta_{n})\leq\theta_{l}\leq\hat{a}(l\Delta_{n})\lor a(l\Delta_{n}),$ such that
$\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}(g)=U_{t}^{n}+\mathcal{R}_{t}^{n},\,\,\,\forall\,0\leq t\leq T,$
@@
$U_{t}^{n}=\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(\theta_{l})\left[\hat{a}(l\Delta_{n})-a(l\Delta_{n})\right],$ (48)
and
@@
Therefore, the conclusion of Theorems 2 and 4 will follow if the processes $U^{n}$ and $\mathcal{R}^{n}$ are asymptotically negligible. Please note that below we will always write $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$, respectively.
@@
Let $g\in\mathfrak{C}_{p,q}^{1}$ with $p,q\geq 1$ and note that in this situation
$\ \mathcal{R}_{t}^{n}=$ $\int_{t}^{m_{n}\Delta_{n}}[g(a(\tau_{n}(s)))-g(a(s))]\mathrm{d}s+\int_{[t/\Delta_{n}]\Delta_{n}}^{t}g(a(\tau_{n}(s)))\mathrm{d}s$ (50)
$-\int_{m_{n}\Delta_{n}}^{\infty}g(a(s))\mathrm{d}s.$
From (11), the mapping $s\mapsto g(a(s))$ is continuous with $\rvert g(a(s))\rvert\leq C_{a(0)}a(s)$. Therefore, uniformly on $t\in[0,T]$
@@
$\int_{t}^{m_{n}\Delta_{n}}\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\mathrm{d}s$ $\lesssim_{a}\int_{0}^{m_{n}\Delta_{n}}[a(\tau_{n}(s))-a(s)]\mathrm{d}s$
$\leq\int_{0}^{\Delta_{n}}[a(0)-a(s)]\mathrm{d}s+\int_{0}^{\Delta_{n}}a(s)\mathrm{d}s$
$-\int_{(N_{n}-1)\Delta_{n}}^{N_{n}\Delta_{n}}a(s)\mathrm{d}s,$
$\lesssim_{a}\Delta_{n}$
This shows that $\mathcal{R}^{n}\overset{u.c.p}{\rightarrow}0$. Finally, since $\rvert\dot{g}(x)\rvert\lesssim\rvert x\rvert^{p}\mathbf{1}_{\rvert x\rvert\leq 1}+\rvert x\rvert^{q}\mathbf{1}_{\rvert x\rvert>1}$, we conclude that
$\sup_{t\in[0,T]}\rvert U_{t}^{n}\rvert\lesssim$ $\Delta_{n}\sum_{l=0}^{m_{n}-1}\left(\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{p\land q+1}+\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{p\lor q+1}\right)$ (51)
@@
$=$ $\mathrm{O}_{\mathbb{P}}(\frac{m_{n}}{n}(n\Delta_{n})^{\frac{1-p\land q}{2}})+\mathrm{O}_{\mathbb{P}}(1/\sqrt{n\Delta_{n}})+\mathrm{o}(1).$
where in the last step we applied Lemma 4 and the fact that $p,q\geq 1$. If $p,q>1$ and $m_{n}=n$ (as in Theorem 2), the right-hand side of (51) is clearly asymptotically negligible. On the other hand if $p\land q=1$ and $m_{n}=N_{n}$ satisfies (12) we have that the right-hand side of (51) is of order $\mathrm{O}_{\mathbb{P}}(\frac{N_{n}}{n})$, which by assumption converges to 0 as $n\rightarrow\infty$.∎
@@
Therefore we only need to show that $B_{n}^{\prime}\overset{\mathbb{P}}{\rightarrow}B$. Now, using the random variables defined in (33), we set
@@
where $\xi_{j,l,n}^{\prime(1)}=\chi_{j}\beta_{j,j-1,l}^{(2)}$ as well as put (recall the decomposition in (32))
@@
Now, a simple application of equation (63) in Appendix A and the definition of $R_{j,l,n}$, let us conclude that $\mathbb{E}(\mathcal{E}_{n}^{\prime})\rightarrow 0.$ This shows, just as above, that the limit of $B_{n}^{\prime\prime}$ is the same as the one of $B_{n}^{\prime}$ (ergo of $B_{n}$). Similar arguments used in the proof of Lemma 7 in A allow us to conclude that
@@
The stated convergence now follows by invoking once again (63) in A.∎
4.3 Proof of Theorems 5 and 6
Similarly as in the previous subsection we fix $T>0$ and choose $n$ large enough such that $(m_{n}-1)\Delta_{n}>T$, where, as before $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$, respectively. Once again we will unify our proofs by means of (46). Specifically, for every function $g$ of class $\mathcal{C}^{2}$ we have that
$Z_{t}^{n}:=\sqrt{n\Delta_{n}}\left(\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}^{n}(g)\right)=V_{t}^{n}+\sum_{\ell=1}^{5}U_{t}^{n,\ell},$
@@ 4.3 Proof of Theorems 5 and 6
$V_{t}^{n}:=\sqrt{n\Delta_{n}}\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))\left[\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\right],$ (52)
and
@@
$U_{t}^{n,2}=$ $\sqrt{n\Delta_{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\frac{(n-l)}{n}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}\dot{g}(a(l\Delta_{n}))[a(s)-a(l\Delta_{n})]\mathrm{d}s;$
$U_{t}^{n,3}=$ $-\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))a(l\Delta_{n})l\Delta_{n};$
$U_{t}^{n,4}=$ $\frac{1}{n}S_{n}\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\dot{g}(a(l\Delta_{n}))(X_{n\Delta_{n}}-X_{l\Delta_{n}});$
@@
The next result shows that within our framework, the leading term is $V^{n}$.
@@
That $U_{t}^{n,1}$ is negligible when $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$(i.e. when $\mathcal{S}^{n}(g)=\Psi^{n}(g)$) can be shown exactly in the same way as in the proof of Theorems 2 and 4 (just replace $m_{n}$ by $T/\Delta_{n}]$ in (51)). Now, if $g\in\mathfrak{C}_{p,q}^{2}$ and either $m_{n}=n$ or $m_{n}=N_{n}$, we deduce as in (51) that uniformly on $t\in[0,T]$
@@
Since $m_{n}/n$ is bounded, we have that $U^{n,1}\overset{u.c.p}{\longrightarrow}0$ when either $p,q>1$ or $p\land q=1$ and $m_{n}=N_{n}$. If instead $0\leq p\land q<1$, suppose that $m_{n}=N_{n}$ and that Assumption 2 holds with $\kappa$ satisfying (16). To check that in this situation the right-hand side of (53) is still negligible as $n\rightarrow\infty$, it is enough to see that $\sqrt{n\Delta_{n}}\frac{N_{n}}{n}\rightarrow 0.$ If $\kappa\leq 1/2$ the result is obvious, so suppose that $1/2<\kappa<\frac{1}{2}(1+\frac{1}{\varpi}).$ Then by noticing that $\frac{1}{1-2(1-\kappa)}>\varpi$, we obtain that

### v3

@@ 4.2. Proof of Theorems 1 - 4
Without loss assume that $g\geq 0$ (otherwise analyze $g^{+}$ and $g^{-}$ separately). Thus, we only need to check that for every $0\leq t\leq T$, $\Psi_{t}^{n}(g)\overset{\mathbb{P}}{\rightarrow}\Psi_{t}(g)$. Suppose first that $g$ is bounded. Then
@@
Theorem 1 in [19] along with the boundedness of $g$ allow us to apply the Dominated Convergence twice to conclude that the right-hand side of the previous inequality goes to 0 as $n\rightarrow\infty$. Suppose now that $g$ is of polynomial growth of order $q>0$ (when $q=0$ $g$ is bounded). Let $\psi:\mathbb{R}^{+}\rightarrow[0,1]$ be a function of class $\mathcal{C}^{\infty}$ such that $\mathbf{1}_{[1,\infty)}\leq\psi\leq\mathbf{1}_{[1/2,\infty)}$, and for every $m\in\mathbb{N}$, set $\psi_{m}(x):=1-\psi(\rvert x\rvert/m)$ as well as $\tilde{\psi}_{m}(x)=\psi(\rvert x\rvert/m).$ Write
@@
where $g_{m}=g\psi_{m}$ and $\tilde{g}_{m}=g\tilde{\psi}_{m}$. Since $g_{m}$ is continuous and bounded, the first part of the proof gives that $\Psi_{t}^{n}(g_{m})\overset{\mathbb{P}}{\rightarrow}\int_{0}^{t}g_{m}(a(s))\mathrm{d}s$. In view that $\int_{0}^{t}g_{m}(a(s))\mathrm{d}s\rightarrow\Psi_{t}(g),$ as $m\rightarrow\infty$, we are left to show that for every $\eta>0$
@@
Since $g$ has polynomial growth, it holds that $\rvert\tilde{g}_{m}(x)\rvert\leq C_{q}\rvert x\rvert^{q}\mathbf{1}_{\rvert x\rvert>m}$. Hence, (43) follows if
@@
From (25)
$Y_{n}^{m}(t)\lesssim_{q}\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert\tilde{a}(l\Delta_{n})\rvert^{q}\mathbf{1}_{\{\rvert\hat{a}(l\Delta_{n})\rvert>m\}}+\rvert\frac{1}{n}S_{n}\rvert^{q}\frac{1}{(n\Delta_{n})^{q}}\Delta_{n}\sum_{l=0}^{[t/\Delta_{n}]-1}\rvert X_{n\Delta_{n}}-X_{l\Delta_{n}}\rvert^{q}.$
Denote by $Y_{n}^{m,1}$ and $Y_{n}^{2}$ the last two terms. As in the proof of Lemma 4, $Y_{n}^{2}\overset{\mathbb{P}}{\rightarrow}0$. Thus,
@@
$\lesssim$ $\frac{t}{(n\Delta_{n})^{q/2}}+\sum_{l=0}^{[t/\Delta_{n}]-1}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}a(s)^{q}\mathrm{d}s\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m).$
By Markov’s inequality
$\mathbb{P}(\rvert\hat{a}(l\Delta_{n})\rvert>m)\leq\frac{1}{m^{q}}\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})\rvert^{q}).$
Using (25) together with Lemma 3, we obtain that $\mathbb{E}(\rvert\tilde{a}(l\Delta_{n})\rvert^{q})$ is bounded uniformly in $n$ and $l$. Therefore,
$\mathbb{E}(\rvert Y_{n}^{m,1}\rvert)\lesssim\frac{1}{(n\Delta_{n})^{q/2}}+\frac{1}{m^{q}}\int_{0}^{t}a(s)^{q}\mathrm{d}s.$
@@
Our next goal is to unify the proof of Theorems 2 and 4. To this end, we write $\mathcal{S}^{n}(g)$ and $\mathcal{S}(g)$ to denote any of the pairs $(\Psi^{n}(g),\Psi(g))$, $(\Lambda^{n}(g),\Lambda(g))$, or $(\bar{\Lambda}^{n}(g),\bar{\Lambda}(g))$. With this notation,
@@
If $g\in\mathcal{C}^{1}$, the Mean Value Theorem yields
$\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}(g)=U_{t}^{n}+\mathcal{R}_{t}^{n},\quad 0\leq t\leq T,$ (48)
@@
$U_{t}^{n}=\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)g^{\prime}(\theta_{l})\left[\hat{a}(l\Delta_{n})-a(l\Delta_{n})\right],$ (49)
for some $\hat{a}(l\Delta_{n})\land a(l\Delta_{n})\leq\theta_{l}\leq\hat{a}(l\Delta_{n})\lor a(l\Delta_{n})$, and
@@
Hence Theorems 2 and 4 follow once the processes $U^{n}$ and $\mathcal{R}^{n}$ are shown to be asymptotically negligible. In what follows we write $m_{n}=n$ or $m_{n}=N_{n}$, depending on whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$, respectively.
@@
Let $g\in\mathfrak{C}_{p,q}^{1}$ with $p,q\geq 1$ and recall the decomposition (48). From the definition of $\mathcal{R}_{t}^{n}$ we get
$\mathcal{R}_{t}^{n}=\int_{t}^{m_{n}\Delta_{n}}[g(a(\tau_{n}(s)))-g(a(s))]\mathrm{d}s+\int_{[t/\Delta_{n}]\Delta_{n}}^{t}g(a(\tau_{n}(s)))\mathrm{d}s-\int_{m_{n}\Delta_{n}}^{\infty}g(a(s))\mathrm{d}s.$ (51)
Since $g(a(\cdot))$ is continuous and satisfies $\rvert g(a(s))\rvert\leq C_{a(0)}a(s)$, (see (11)), we have that, uniformly on $t\in[0,T]$,
@@
$\int_{t}^{m_{n}\Delta_{n}}\rvert g(a(\tau_{n}(s)))-g(a(s))\rvert\mathrm{d}s\lesssim_{a}\int_{0}^{m_{n}\Delta_{n}}[a(\tau_{n}(s))-a(s)]\mathrm{d}s\lesssim_{a}\Delta_{n},$
where in the last inequality we further used (41). Hence $\mathcal{R}^{n}\overset{u.c.p}{\rightarrow}0$. Let us now turn our attention to $U^{n}$. Since $\rvert g^{\prime}(x)\rvert\lesssim\rvert x\rvert^{p}\mathbf{1}_{\rvert x\rvert\leq 1}+\rvert x\rvert^{q}\mathbf{1}_{\rvert x\rvert>1}$, we obtain
$\sup_{t\in[0,T]}\rvert U_{t}^{n}\rvert\lesssim$ $\Delta_{n}\sum_{l=0}^{m_{n}-1}\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{p\land q+1}+\Delta_{n}\sum_{l=0}^{m_{n}-1}\rvert\hat{a}(l\Delta_{n})-a(l\Delta_{n})\rvert^{p\lor q+1}$ (52)
@@
Applying Lemma 4 yields
$\sup_{t\in[0,T]}|U_{t}^{n}|=\mathrm{O}_{\mathbb{P}}(\frac{m_{n}}{n}(n\Delta_{n})^{\frac{1-p\land q}{2}})+\mathrm{O}_{\mathbb{P}}(1/\sqrt{n\Delta_{n}})+\mathrm{o}(1).$
If $p,q>1$ and $m_{n}=n$ (as in Theorem 2), the right-hand side of the previous estimate converges to zero. If $p\land q=1$ and $m_{n}=N_{n}$ satisfies (12) (as in Theorem 4), the bound in (52) becomes $\mathrm{O}_{\mathbb{P}}(\frac{N_{n}}{n})$, which also converges to zero by assumption. Thus $U^{n}$ and $\mathcal{R}^{n}$ are asymptotically negligible, completing the proof.∎
@@
Therefore, it suffices to prove that $B_{n}^{\prime}\overset{\mathbb{P}}{\rightarrow}B$. Now, using the random variables introduced in (33) and (32), we set
@@
where $\xi_{j,l,n}^{\prime(1)}=\chi_{j}\beta_{j,j-1,l}^{(2)}$, and let
@@
By (32), (63) in Appendix A, and the definition of $R_{j,l,n}$, we get that $\mathbb{E}(\mathcal{E}_{n}^{\prime})\rightarrow 0$. This shows, just as above, that $B_{n}^{\prime\prime}$ and $B_{n}^{\prime}$ have the limit (and therefore the same limit as $B_{n}$). Arguing as in Lemma 7 in Appendix A, we conclude that
@@
The stated convergence now follows by invoking once again (63) in Appendix A.∎
4.3. Proof of Theorems 5 and 6
We begin by outlining the proof of the central limit theorems. Starting from a decomposition of the estimation error into a leading term and several remainder terms, we first show that the remainder terms are asymptotically negligible. The leading term is then rewritten as a sum of martingale differences plus a negligible error term. The proof proceeds by establishing tightness, reducing the convergence of the finite-dimensional to a martingale central limit theorem via the Cramér–Wold device, and verifying the convergence of the conditional variance and the Lyapunov condition.
Similarly to the previous subsection, fix $T>0$ and choose $n$ large enough such that $(m_{n}-1)\Delta_{n}>T$, where, as before $m_{n}=n$ or $m_{n}=N_{n}$, depending whether $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq n-1}$ or $c_{l}^{n}(t)=\mathbf{1}_{[t/\Delta_{n}]\leq l\leq N_{n}-1}$, respectively.
Once again, we unify the proofs by means of (46). Specifically, for $g\in\mathcal{C}^{2}$, define
$Z_{t}^{n}:=\sqrt{n\Delta_{n}}\left(\mathcal{S}_{t}^{n}(g)-\mathcal{S}_{t}(g)\right).$
Then
$Z_{t}^{n}=V_{t}^{n}+\sum_{\ell=1}^{5}U_{t}^{n,\ell},$
@@ 4.3. Proof of Theorems 5 and 6
$V_{t}^{n}:=\sqrt{n\Delta_{n}}\Delta_{n}\sum_{l=0}^{n-1}c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))\left[\tilde{a}(l\Delta_{n})-\mathbb{E}(\tilde{a}(l\Delta_{n}))\right],$ (53)
is the leading term, and the remainder terms are given by
@@
$U_{t}^{n,2}=$ $\sqrt{n\Delta_{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)\frac{(n-l)}{n}\int_{l\Delta_{n}}^{(l+1)\Delta_{n}}g^{\prime}(a(l\Delta_{n}))[a(s)-a(l\Delta_{n})]\mathrm{d}s;$
$U_{t}^{n,3}=$ $-\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))a(l\Delta_{n})l\Delta_{n};$
$U_{t}^{n,4}=$ $\frac{1}{n}S_{n}\sqrt{\frac{\Delta_{n}}{n}}\sum_{l=0}^{n-1}c_{l}^{n}(t)g^{\prime}(a(l\Delta_{n}))(X_{n\Delta_{n}}-X_{l\Delta_{n}});$
@@
The next lemma shows that, under our assumptions, all remainder terms are asymptotically negligible and the asymptotic behaviour is therefore determined by $V^{n}$.
@@
We treat the terms $U^{n,i}$ separately.
Case $i=1$. That $U_{t}^{n,1}$ is negligible when $c_{l}^{n}(t)=\mathbf{1}_{0\leq l\leq[t/\Delta_{n}]-1}$ follows exactly as in the proof of Theorems 2 and 4 by replacing $m_{n}$ with $T/\Delta_{n}]$ in (52). Now, if $g\in\mathfrak{C}_{p,q}^{2}$ and either $m_{n}=n$ or $m_{n}=N_{n}$, the same argument yields that, uniformly on $t\in[0,T]$,
@@
Since $m_{n}/n$ is bounded, we have that $U^{n,1}\overset{u.c.p}{\longrightarrow}0$ whenever $p,q>1$, or $p\land q=1$ and $m_{n}=N_{n}$. When $0\leq p\land q<1$, Assumption 2 together with (16) implies that $\sqrt{n\Delta_{n}}\frac{N_{n}}{n}\rightarrow 0$. Indeed, if $\kappa\leq 1/2$ the claim is obvious. If $1/2<\kappa<\frac{1}{2}(1+\frac{1}{\varpi})$, we obtain that